Can someone please help solve
1/5x+py =x
4/5x +(1-p)=y
x+y=1
solve for x,y
x=
y=

Answer by RAY100(1637)

1) .2x+py=x
2).8x+1-p=y
3)x+y=1
from (1) .2x +py =x, py=x-.2x=.8x, py=.8x
from(2) .8x +1-p =y, subst from (1), py +1-p=y, py-p=y-1, p(y-1)=(y-1), p=1 answer
but py=.8x, then y=.8x
subst in (3) x+y=1, x+.8x=1, 1.8x=1, x=1/1.8 , x=1/(1 4/5)=1/(9/5), x=5/9 answer
from (3) x+y=1, 5/9 +y =1, y=1-5/9, y=4/9 answer
checking
(3) 5/9 +4/9 +1 ok
(1) 1/5*5/9 +1*4/9 = 5/9 , 1/9 + 4/9=5/9 ok
(2) 4/5 *5/9 +1-1 = 4/9 , 4/9 = 4/9 ok